### v2

# Distributed Certifiably Correct Range-Aided SLAM

Abstract
Reliable simultaneous localization and mapping (SLAM) algorithms are necessary for safety-critical autonomous navigation. In the communication-constrained multi-agent setting, navigation systems increasingly use point-to-point range sensors as they afford measurements with low bandwidth requirements and known data association. The state estimation problem for these systems takes the form of range-aided (RA) SLAM. However, distributed algorithms for solving the RA-SLAM problem lack formal guarantees on the quality of the returned estimate. To this end, we present the first distributed algorithm for RA-SLAM that can efficiently recover certifiably globally optimal solutions. Our algorithm, distributed certifiably correct RA-SLAM (DCORA), achieves this via the Riemannian Staircase method, where computational procedures developed for distributed certifiably correct pose graph optimization are generalized to the RA-SLAM problem. We demonstrate DCORA’s efficacy on real-world multi-agent datasets by achieving absolute trajectory errors comparable to those of a state-of-the-art centralized certifiably correct RA-SLAM algorithm. Additionally, we perform a parametric study on the structure of the RA-SLAM problem using synthetic data, revealing how common parameters affect DCORA’s performance.
I Introduction
Realizing autonomy as a ubiquitous technology requires formal guarantees on performance. This is increasingly true for multi-agent autonomy, where coordinated planning, control, and perception must remain safety-critical for cyber-physical systems [1] and post-disaster evaluations [2], among other applications. Parallel to the long-standing guarantees established by the planning [3] and control [4, 5] communities, a concerted effort over the past decade has been directed toward endowing perception with formal guarantees, where the prevailing algorithms underlying these methods are certifiably correct (i.e., algorithms that recover certifiably globally optimal solutions of generally-intractable problems within a restricted but operationally relevant set of problem instances [6, 7]).
In this work, we advance certifiable perception for multi-agent autonomy by presenting DCORA, the first certifiably correct algorithm for distributed range-aided simultaneous localization and mapping (RA-SLAM). The RA-SLAM problem estimates a set of pose and landmark locations from pairwise relative pose and range measurements, and DCORA solves this problem while obeying the topology of an underlying distributed communication graph (Fig. 1).
Figure 1: (a) Graphical representation of the RA-SLAM problem in the distributed setting. In DCORA, synchronized agents (e.g., $\mathcal{A}_{\alpha},\mathcal{A}_{\beta},\mathcal{A}_{\gamma}$) jointly estimate their state in a common frame by performing local computation steps and exchanging information over the distributed communication graph shown in (b). States, including poses (circles) and landmarks (stars), are interconnected via relative pose (black lines) and range (blue lines) measurements. To facilitate efficient communication and computation, states are labeled private (black) or public (red), while measurements exist as inter-agent loop closures (dotted lines), intra-agent loop closures (dash-dotted lines), or odometry (solid lines) as discussed in Section III-C.
Specifically, DCORA applies distributed Riemannian optimization techniques [8] to solve the (convex) semidefinite programming (SDP) relaxation of a maximum likelihood (ML) formulation of RA-SLAM [9]. In turn, solving the SDP allows DCORA to recover globally optimal RA-SLAM solutions within practically relevant noise regimes. We solve the SDP via the Riemannian Staircase method [10, 11], for which we use the Riemannian block coordinate descent (RBCD) algorithm [8] to identify first-order critical points (i.e., points where the Riemannian gradient of the objective function vanishes). We leverage established Karush-Kuhn-Tucker (KKT)-based conditions (which require a certain dual certificate matrix, when evaluated at the global minima, to be positive semidefinite) of the RA-SLAM problem [12, 9] to verify global optimality, and apply distributed saddle escape techniques to descend from suboptimal critical points (if necessary). We benchmark DCORA against a state-of-the-art centralized certifiably correct RA-SLAM algorithm on real-world datasets, demonstrating comparable absolute trajectory error (ATE) performance. Furthermore, we perform parameter sweeps across synthetic datasets to observe how common parameters (namely, the number of agents, the number of landmarks, and range measurement density) in RA-SLAM problems affect DCORA’s performance in terms of convergence rate, solution precision, and SDP tightness. These contributions are summarized as follows:
  1. 1.
    The first distributed RA-SLAM algorithm capable of producing certifiably globally optimal estimates.
  2. 2.
    A benchmark comparison of our algorithm against a state-of-the-art centralized certifiably correct RA-SLAM algorithm on real-world datasets.
  3. 3.
    A parametric study on DCORA’s convergence rate, solution quality, and SDP tightness using synthetic datasets.
  4. 4.
    An open-source implementation of our algorithm [^1] https://github.com/adthoms/dcora.
II Related Works
We briefly review works related to DCORA’s method class, distributed Riemannian optimization algorithms amenable to said method class, and existing distributed RA-SLAM algorithms.
II-A Certifiably Correct Perception via the Riemannian Staircase
DCORA depends on a formulation of the RA-SLAM problem that allows an SDP relaxation to be constructed and subsequently solved through a series of rank-restricted SDPs of considerably lower dimension than the original SDP. Furthermore, these specific rank-restricted SDPs are equivalent to optimization over the Stiefel and Oblique manifolds, leading DCORA to instead solve a series of (smooth) manifold optimization problems. This approach – solving an SDP via a series of lower-dimensional manifold optimization problems – is known as the Riemannian Staircase method [10, 11].
The first work to apply the Riemannian Staircase for certifiable perception was that of [13], which presented SE-Sync, a certifiably correct algorithm for pose graph optimization (PGO). However, SE-Sync is an inherently centralized algorithm, which presents challenges in communication-constrained multi-agent settings. Tian et al. [8] addressed this with the DC2-PGO algorithm, which (to the best of our knowledge) is the only distributed certifiably correct algorithm for robotic perception. DC2-PGO presented distributed algorithms for the two key procedures of SE-Sync: optimization and certification. The capabilities of certifiable perception were then further pushed in our previous work [9], which developed a novel SDP formulation for RA-SLAM and presented CORA, a centralized certifiably correct algorithm for RA-SLAM via the Riemannian Staircase.
II-B Distributed Riemannian Optimization Algorithms
Knuth and Barooah [14, 15, 16] presented the first distributed Riemannian optimization algorithm for robotic perception, called D-RPGO. The author’s method poses PGO as a product of Stiefel manifolds, which is solved via a provably correct explicit gradient descent law. D-RPGO was developed in parallel to distributed Riemannian optimization algorithms developed for computer vision [17, 18], which are also based on gradient descent. Tian et al. [8] developed the RBCD algorithm (and an accelerated variant RBCD++) as the core optimization procedure of DC2-PGO. RBCD/RBCD++ performs distributed block-coordinate descent over the product of (arbitrary smooth) Riemannian manifolds, and provably converges to first-order critical points with a global sublinear rate. These methods assume synchronization among agents, and Tian et al. [19] addressed this limitation by developing ASAPP, the first asynchronous and provably convergent algorithm for distributed PGO. Importantly, the authors show how ASAPP generalizes to settings where the feasible set is the product of Riemannian manifolds. Distributed Riemannian optimization for PGO was further advanced by Li et al. [20], who proposed combined multilevel graph partitioning (for optimization load balancing) with an accelerated Riemannian optimization method, called IRBCD. IRBCD enjoys the same convergence properties as RBCD/RBCD++ while demonstrating improvements to solution quality and convergence rate on benchmark datasets. Lastly, Fan and Murphey [21] developed a majorization-minimization algorithm for distributed PGO that is guaranteed to converge to first-order critical points and leverages Nesterov acceleration and adaptive restart (similar to RBCD++) for accelerated convergence. We note that distributed Riemannian optimization algorithms have been developed for active robotic perception [22] and are theoretically compatible with DCORA’s method class.
II-C Distributed RA-SLAM Algorithms
Distributed SLAM algorithms have been developed to solve specific instances of the RA-SLAM problem. Such algorithms have been developed to fuse range measurements between robots [23, 24, 25, 26, 27] and into compound relative pose estimates between robots [28, 29, 30, 31, 32]; however, none provide a unified approach across all state and measurement combinations that arise in RA-SLAM. Moreover, these methods provide no guarantees (a priori or a posteriori) on the quality of the returned estimate. We note that several distributed back-end solvers [33, 34, 35, 36, 37, 38, 39] are capable of solving the RA-SLAM problem; however, they too lack formal guarantees on solution quality. That is, these methods rely on local search techniques (e.g., Gauss-Newton [40]) and are thus susceptible to significantly suboptimal critical points. Even worse, certain local search techniques (e.g., Gaussian belief propagation [41]) lack convergence guarantees. The limitations of existing distributed RA-SLAM algorithms and distributed back-end solvers thus necessitate a new approach for safety-critical navigation.
III Methodology
In this section, we first formulate the RA-SLAM problem and present its connection to Riemannian optimization. Next, we present the DCORA algorithm with a discussion on enabling efficient communication and computation via variable ownership.
III-A RA-SLAM Problem Formulation and the Riemannian Staircase
To solve the RA-SLAM problem via the Riemannian Staircase, we begin with the following ML formulation of RA-SLAM
Problem 1 (ML formulation of RA-SLAM).
Given relative pose measurements $\{\tilde{R}_{ij},\tilde{t}_{ij}\}_{(i,j)\in\mathcal{E}_{p}}$ and point-to-point distance measurements $\{\tilde{r}_{ij}\in\mathbb{R}\}_{(i,j)\in\mathcal{E}_{r}}$, solve
$\min\limits_{\begin{subarray}{c}R_{i}\in\operatorname{{\operatorname{SO}}(d)}\\ t_{i}\in\mathbb{R}^{d}\end{subarray}}{}{}$ $\sum\limits_{(i,j)\in\mathcal{E}_{p}}\kappa_{ij}\lVert R_{j}-R_{i}\tilde{R}_{ij}\rVert_{F}^{2}$ (1)
$+$ $\sum\limits_{(i,j)\in\mathcal{E}_{p}}\tau_{ij}\left\lVert t_{j}-t_{i}-R_{i}\tilde{t}_{ij}\right\rVert_{2}^{2},$
$+$ $\sum_{(i,j)\in\mathcal{E}_{r}}{\rho_{ij}}(\lVert t_{j}-t_{i}\rVert_{2}-\tilde{r}_{ij})^{2}$
where $d$ is the dimension of the problem (e.g., 2D or 3D), $\|\cdot\|_{F}$ denotes the matrix Frobenius norm, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\mathcal{E}_{p}$ and $\mathcal{E}_{r}$ are index sets for the relative pose measurements and range measurements (respectively), and $\kappa_{ij},\tau_{ij},\rho_{ij}\in\mathbb{R}_{+}$ are measurement precisions for the relative pose and range measurements (respectively).
Without loss of generality, 1 does not distinguish landmark variables from poses (as discussed in [9, Appendix E]) to simplify mathematical presentation. To recover globally optimal solutions to 1, [9] developed a (non-convex) quadratically constrained quadratic programming (QCQP) relaxation of 1, which enabled a (convex) semidefinite relaxation to be constructed via Shor’s relaxation [42]. The QCQP relaxation is as follows
Problem 2 (QCQP relaxation of RA-SLAM).
Given the measurements of 1, find the variables, $\{R_{i},t_{i}\}_{i=1}^{n}\cup\{r_{ij}\}_{(i,j)\in\mathcal{E}_{r}}$ , which solve
$\min\limits_{\begin{subarray}{c}R_{i}\in\mathbb{R}^{d\times d}\\ t_{i}\in\mathbb{R}^{d}\\ r_{ij}\in\mathbb{R}^{d}\end{subarray}}{}{}$ $\sum\limits_{(i,j)\in\mathcal{E}_{p}}\kappa_{ij}\lVert R_{j}-R_{i}\tilde{R}_{ij}\rVert_{F}^{2}$ (2)
$+$ $\sum\limits_{(i,j)\in\mathcal{E}_{p}}\tau_{ij}\left\lVert t_{j}-t_{i}-R_{i}\tilde{t}_{ij}\right\rVert_{2}^{2}$
$+$ $\sum_{(i,j)\in\mathcal{E}_{r}}{\rho_{ij}}\lVert t_{j}-t_{i}-\tilde{r}_{ij}r_{ij}\rVert_{2}^{2}$
subject to $R_{i}^{\top}R_{i}=I_{d},~{}i=1,\ldots,n$
$\lVert r_{ij}\rVert_{2}^{2}=1,~{}\forall(i,j)\in\mathcal{E}_{r}~{}.$
2 relaxes the special orthogonal constraint $R_{i}\in\operatorname{{\operatorname{SO}}(d)}$ to an orthogonality constraint, $R_{i}\in\operatorname{{\operatorname{O}}(d)}$, and introduces auxiliary unit-norm vector variables, $r_{ij}\in\mathbb{R}^{d}$, as a reformulation of the range cost terms. From 2, an SDP relaxation can be constructed via Shor’s relaxation [42]
Problem 3 (Semidefinite relaxation of RA-SLAM).
Find $Z\in\mathbb{R}^{k\times k}$ that solves
$\min_{Z\in\mathbb{S}^{k}_{+}}$ $\operatorname{tr}(QZ)$ (3)
subject to $\operatorname{tr}(A_{i}Z)=b_{i},\quad\forall i=1,\dotsc,m,$
where $k\triangleq n(d+1)+l$, $l\triangleq|\mathcal{E}_{r}|$ is the number of range measurements, and $Q$ and $A_{i}$ are defined in [9, Appendix B] and [9, Appendix C], respectively.
The exact steps to form this relaxation are outlined in [9, Appendix D]. Key to our approach is recognition that 3 should often admit low-rank solutions (i.e., solutions $Z^{\ast}$ for which $\operatorname{rank}Z^{\ast}\ll k$). Indeed, this low-rank behavior has been found in practice [9]. Thus, instead of solving 3 directly via standard SDP solvers, one may instead leverage the low-rank structure by optimizing over a low-dimensional factorization $X^{\top}X=Z,X\in\mathbb{R}^{p\times k}$ for user-selectable parameter $p\ll k$. This technique, known as the Burer-Monteiro method [43, 44, 45], arrives at the following problem
Problem 4 (Rank-restricted SDP).
$\min_{X\in\mathbb{R}^{p\times k}}$ $\operatorname{tr}(QX^{\top}X)$ (4)
subject to $\operatorname{tr}(A_{i}X^{\top}X)=b_{i},\quad\forall i=1,\dotsc,m,$
in which the factorization implicitly constrains the rank of the solution to be less than or equal to $p$. For our specific problem, $X$ can be structured as
$X\in\mathbb{R}^{p\times k}\triangleq\left[R_{1}\ \dotsb\ R_{n}\mid r_{1}\ \dotsb\ r_{l}\mid t_{1}\ \dotsb\ t_{n}\right].$ (5)
For higher problem dimensions ($p>d$), variables (e.g., $R_{1}$) in (5) can be thought of as higher-dimensional generalizations of the original variables of the QCQP (2); 4 thus represents a higher-dimensional relaxation of 2. In fact, analysis of the constraints of 4 reveals that the problem can be equivalently posed as Riemannian manifold optimization [9]
Problem 5 (Problem 4’s formulation as Riemannian manifold optimization).
Given the measurements of Problem 1, find the variables, $\{R_{i},t_{i}\}_{i=1}^{n}\cup\{r_{ij}\}_{(i,j)\in\mathcal{E}_{r}}$ , which solve
$\min\limits_{\begin{subarray}{c}R_{i}\in\operatorname{St}(p,d)\\ t_{i}\in\mathbb{R}^{p}\\ r_{ij}\in S^{p-1}\end{subarray}}{}{}$ $\sum\limits_{(i,j)\in\mathcal{E}_{p}}\kappa_{ij}\lVert R_{j}-R_{i}\tilde{R}_{ij}\rVert_{F}^{2}$ (6)
$+$ $\sum\limits_{(i,j)\in\mathcal{E}_{p}}\tau_{ij}\left\lVert t_{j}-t_{i}-R_{i}\tilde{t}_{ij}\right\rVert_{2}^{2},$
$+$ $\sum_{(i,j)\in\mathcal{E}_{r}}{\rho_{ij}}\left\lVert t_{j}-t_{i}-\tilde{r}_{ij}r_{ij}\right\rVert_{2}^{2}$
where $p$ is the rank of the problem (i.e., $p\geq d$), $\operatorname{St}(p,d)$ is the Stiefel manifold of orthonormal $d$-frames in $p$-dimensional space [46, Ch. 7.3], and $S^{p-1}$ is the unit-sphere in $\mathbb{R}^{p}$.
This manifold optimization representation (5) directly generalizes the original QCQP (2), as the Stiefel manifold $\operatorname{St}(p,d)$ generalizes the orthogonal group $\operatorname{{\operatorname{O}}(d)}$ and the auxiliary variables $r_{ij}$ are still constrained to be unit-norm but can be of arbitrary dimension $p$. 5 is directly optimized in the DCORA algorithm. For notational convenience, we refer to the product manifold defined by 5’s variables as $\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p,n,l)\triangleq\operatorname{St}(p,d)^{n}\times(S^{p-1})^{l}\times\mathbb{R}^{p\times n}\subset\mathbb{R}^{p\times k}$.
III-B DCORA Algorithm
We briefly summarize the DCORA algorithm (Algorithm 1), which is a generalization of the DC2-PGO algorithm [8] to the RA-SLAM problem.
Algorithm 1 Distributed Certifiably Correct Range-Aided SLAM (DCORA)
0:  Initial rank $p_{0}\geq d$ for the Riemannian Staircase.
0:  A feasible solution $T\in\operatorname{{\operatorname{SE}}(d)}^{n}$ to 1 and the lower bound $f^{\ast}_{\text{SDP}}$ on 1’s optimal value.
1:  Obtain initial point $X_{\text{init}}\in\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p,n,l)$ for $p=d$ via distributed initialization.
2:  Lift initial point to initial rank $X=Y_{\text{rand}}X_{\text{init}}$.
3:  for $p=p_{0},p_{0}+1,\dots$ do
4:     Compute first-order critical point of 4 [8, Algorithm 3] $X^{\ast}\leftarrow\text{RBCD}(X)$.
[…]
6:     Construct corresponding dual certificate matrix [9, Theorem 1] $S(X^{\ast})\triangleq Q+\sum_{i=1}^{m}\lambda_{i}A_{i}$, where $\lambda\in\mathbb{R}^{m}\triangleq\left[\lambda_{1},\ldots,\lambda_{m}\right]$ are the Lagrange multipliers at KKT point $X^{\ast}$.
7:     Compute minimum eigenpair [8, Algorithm 6] $(\lambda,v)\leftarrow\text{MinEig}(S(X^{\ast}))$.
8:     if $\lambda\geq 0$ then
9:        return  $X^{\ast}$.
10:     else
11:        Construct second-order descent direction $\dot{X}_{+}\triangleq[0\ v^{\top}]^{\top}$, where $v\in\mathbb{R}^{k}$ is the eigenvector corresponding to the smallest eigenvalue of $S(X^{\ast})$ which satisfies $v^{\top}S(X^{\ast})v<0$.
12:        Descend from $X^{\ast}$ [8, Algorithm 7] $X\leftarrow\text{EscapeSaddle}(X^{\ast},\dot{X}_{+})$.
13:     end if
14:  end for
15:  Recover the optimal value of the SDP relaxation $f^{\ast}_{\text{SDP}}=\langle Q,{X^{\ast}}^{\top}X^{\ast}\rangle_{F}$, where $\langle\cdot,\cdot\rangle_{F}$ is the Frobenius inner product.
16:  Project $X^{\ast}$ to the feasible set of 1 [9, Algorithm 3] via distributed rounding to obtain $T\in\operatorname{{\operatorname{SE}}(d)}^{n}$.
17:  return  $T,f^{\ast}_{\text{SDP}}$.
An initial point $X_{\text{init}}\in\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p=d,n,l)$ is obtained via distributed initialization. This initial point is lifted to rank $p_{0}$: $X=Y_{\text{rand}}X_{\text{init}}$, where $Y_{\text{rand}}\in\operatorname{St}(p_{0},d)$ is a random Stiefel element sampled from the initial rank $p_{0}\geq d$ of the Riemannian staircase.
From this lifted initial point, the Riemannian Staircase methodology is applied. First, 5 is optimized over via RBCD. Once RBCD has returned a first-order critical point $X^{\ast}$, $X^{\ast}$ is lifted to the next rank $p+1$ in the Riemannian staircase, where a distributed verification step either certifies $X^{\ast}$ as globally optimal or constructs a second-order descent direction to descend from $X^{\ast}$. If $X^{\ast}$ is not globally optimal, then saddle point escape is performed via line search along the computed descent direction and RBCD is used to optimize from the resulting point. The distributed Riemannian Staircase algorithm continues until a certified globally optimal value $f^{\ast}_{\text{SDP}}$ of 3 is recovered, which is a lower bound on 1’s optimal value. If the SDP is tight, as it often is in practice [9], then the solution is globally optimal for 1. Completing the DCORA algorithm, $X^{\ast}$ is projected to the feasible set of 1 via distributed rounding to recover pose and landmark estimates $T\in\operatorname{{\operatorname{SE}}(d)}^{n}\triangleq\{R_{i}\in\operatorname{{\operatorname{SO}}(d)},t_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$.
III-C DCORA’s Variable Ownership Structure
III-C1 Variable Partitioning
A key notion in DCORA is that of variable ownership. Each decision variable in 5 is owned by a single agent, and variables are partitioned by agent ownership to satisfy RBCD’s input requirement. Specifically, RBCD requires a block partitioning of $X$, where a block $X_{b}\subseteq X$ is associated with the agent responsible for estimating those states. In DC2-PGO, agents naturally take ownership of pose variables that describe their trajectory. With the introduction of external landmarks and range measurements, variable partitioning becomes less obvious for the RA-SLAM problem. As such, we maintain DC2-PGO’s pose ownership rule in DCORA and assume agents own landmarks according to a consensus step [47] and auxiliary unit-norm vector variables corresponding to range measurements collected by their onboard sensors.
III-C2 Variable Labeling
In an agent’s local search step, RBCD requires agents to maintain certain local copies of variables belonging to other agents. To facilitate efficient sharing of these copies, poses and landmarks are labeled public or private according to the following definition
Definition 1.
States that connect to the states of other agents (via pairwise measurements) are called public states. All other states are private.
Thus, agents sharing a pairwise measurement (referred to as an inter-agent loop closure) need only cache the neighboring agent’s public state before the local search step. If the pairwise measurement is a range measurement taken by the other agent, the corresponding auxiliary unit-norm vector variable is also cached. Information related to an agent’s private states, intra-agent loop closures (which connect non-consecutive states owned by a single agent), and odometry remains private.
IV Experiments
Experiments conducted to evaluate DCORA are two-fold. First, we benchmark DCORA against its centralized counterpart, CORA, on public real-world multi-agent RA-SLAM datasets. Comparison is limited to CORA as (to our knowledge) this is the only other certifiably correct algorithm that can solve the general RA-SLAM problem. We refer the interested reader to [9] for an in-depth comparison of CORA against the GTSAM Levenberg–Marquardt optimizer [48], a popular local search back-end solver, to motivate our comparison. Second, multi-agent RA-SLAM datasets are generated by modifying public synthetic PGO datasets. These datasets are modified to assign pose variables to multiple agents, include landmarks, and establish inter- and intra-agent range measurement loop closures with controlled density. Parameter sweeps are performed based on (i) the number of agents, (ii) the number of landmarks, and (iii) range measurement density to study the RA-SLAM problem structure. DCORA’s implementation for evaluation precedes the discussion of these experiments.
IV-A Implementation for Evaluation
Our implementation of DCORA centralizes two procedures from Algorithm 1 for convenience in conducting our experiments. Our experiments do not evaluate DCORA’s runtime performance, and our centralized implementations only affect runtime while not affecting the performance metrics we measure. First, we employ a centralized initialization procedure in line 1 that provides quality initial estimates. Our centralized initialization procedure improves runtime over a distributed implementation by assuming all measurements and states belong to a single agent. This reduces the number of computation steps and avoids inter-agent communication when solving for $X_{\text{init}}$. While the solutions returned by DCORA are initialization independent, quality initial estimates (which may be equally obtained using a distributed implementation) further improve runtime by avoiding the need to increase rank multiple times in the Riemannian Staircase to escape from spurious local minima. Second, we perform centralized minimum eigenpair calculations (as part of the solution verification step in line 7) using the centralized Lanczos-based method described in [49]. This method has the same asymptotic convergence rate as DCORA’s (and DC2-PGO’s) distributed minimum eigenpair calculation, thus serving as a good proxy for the runtime performance of distributed solution verification.
IV-B Real-World Efficacy: DCORA vs. CORA
IV-B1 Datasets
Real-world multi-agent RA-SLAM datasets, including TIERS [50] and UTIAS MR.CLAM [51], are used to benchmark DCORA against CORA. To our knowledge, these are the only real-world multi-agent RA-SLAM datasets publicly available in the literature. TIERS comprises five ground robots with stereo cameras for visual odometry and ultra-wideband sensors to perform inter-agent ranging. One robot was stationary, acting as a landmark, while the other four moved in non-overlapping geometric patterns. MR.CLAM comprises five ground robots and fifteen landmarks. All robots and landmarks were equipped with barcode fiducials, and robots were equipped with cameras to obtain range (and bearing [^2] Bearing measurements have been ignored from MR.CLAM to extract RA-SLAM problem instances.) measurements to the fiducials. The robots obtained odometry from open-loop wheel velocity commands. Sessions 2, 4, 6, and 7 from MR.CLAM are used for evaluation as the remaining sessions had substantial dropouts in ground truth data and are therefore unusable.
IV-B2 Evaluation and Results
[…]
TABLE I: DCORA vs. CORA on Public Real-World Multi-Agent RA-SLAM Datasets
[…]
$f^{\ast}_{\text{SDP}}$ $\epsilon$ Trans. [m] Rot. [$\circ$]
TIERS DCORA 1920.4 1e-2 0.04 0.05
CORA 1921.9 1e-3 0.04 0.05
MR.CLAM 2 DCORA 7032.9 5e-3 0.15 0.12
CORA 7034.6 1e-3 0.17 0.12
MR.CLAM 4 DCORA 4173.2 5e-3 0.13 0.10
CORA 4173.1 1e-3 0.13 0.10
MR.CLAM 6 DCORA 3124.8 5e-3 0.27 0.14
CORA 3124.5 1e-3 0.26 0.14
MR.CLAM 7 DCORA 3091.4 5e-2 0.31 0.14
CORA 3038.4 1e-2 0.27 0.13
  • $\dagger$
    Datasets include TIERS[50] and sessions 2, 4, 6, and 7 of MR.CLAM[51].
  • $\ddagger$
    All solutions are certified.
  • $\ast$
    Bold values indicate best performance.
IV-B3 Discussion
Results demonstrate comparable ATE performance between DCORA and CORA. This is expected, as these algorithms are (mainly) differentiated by the order of their optimization procedures (i.e., first- vs. second-order). The order of optimization does not affect the optimality of the SDP; thus, the certified solution (mathematically) should remain identical. However, RBCD has a sublinear convergence rate, meaning its estimates are not as precise as those provided by CORA’s second-order trust-region method. As a result, DCORA requires a looser certification tolerance to verify optimality, which accounts for the small differences in $f^{\ast}_{\text{SDP}}$ and ATE.
IV-C RA-SLAM Problem Structure in the Distributed Setting
IV-C1 Datasets
To study the RA-SLAM problem structure in the distributed setting, two public synthetic PGO datasets, smallGrid3D[8] and sphere2500[55], are modified (Fig. 2) to generate 24 multi-agent RA-SLAM problem instances.
Figure 2: Modified (a) smallGrid3D[8] and (b) sphere2500[55] pose graph optimization (PGO) datasets used to study the range-aided SLAM problem structure in the distributed setting. Original PGO datasets (with their trajectories illustrated in red) have been modified to generate 24 new datasets, which include multiple agents (not shown), landmarks (blue and green circles), and range measurements (not shown).
Modified datasets evenly and contiguously distribute pose variables among two, four, or eight agents, with each dataset assuming a fixed configuration of eight or 16 landmarks. Range measurements are generated assuming a precision of $\rho_{ij}=100\frac{1}{\text{m}^{2}}$ and a sensing horizon of $r_{\text{max}}=100$m. Range measurements are generated with a probability of 0.5 or 1.0 from each agent’s pose variables to (i) a random agent’s pose and (ii) a random landmark within a distance of $r_{\text{max}}$. We assume landmarks are owned by the agent with the most connecting range measurements, thus minimizing inter-agent communication.
IV-C2 Evaluation and Results
For evaluation, we calculate the relative suboptimality upper bound $(f(X^{k})-f^{\ast}_{\text{SDP}})/f^{\ast}_{\text{SDP}}$ (normalized to account for differences in scale between objective values of the modified PGO datasets) and the Riemannian gradient norm $\|\operatorname{grad}f(X^{k})\|$ at each RBCD iterate $X^{k}$ for rank $p=3$, where $f(X^{k})=\langle Q,(X^{k})^{\top}X^{k}\rangle_{F}$. The former of these measures the relative suboptimality of the rounded estimates recovered from the convex relaxation, and so effectively measures how tight our relaxation is. The latter of these measures how precisely RBCD is able to estimate the critical points recovered from the Riemannian rank-restricted SDPs. Fig. 3 illustrates the evolution of these metrics for the generated multi-agent RA-SLAM problem instances.
Figure 3: Evolution of the normalized relative suboptimality upper bound and the Riemannian gradient norm at each RBCD iterate for rank three in the Riemannian staircase. Modified datasets (a) smallGrid3D[8] and (b) sphere2500[55] have the following parameters: (i) landmarks: 8, range measurement probability: 0.5; (ii) landmarks: 16, range measurement probability: 0.5; (iii) landmarks: 8, range measurement probability: 1.0; (iv) landmarks: 16, range measurement probability: 1.0; for two-agent (blue), four-agent (orange), and eight-agent (green) systems. All plots are log-linear.
IV-C3 Discussion
As observed in Fig. 3, convergence rate and solution precision are predominantly affected by the number of agents. This trend is most clear on the modified smallGrid3D dataset, where fewer RBCD iterations are required to converge to a lower Riemannian gradient norm, thus suggesting faster convergence to more precisely estimated critical points. For the modified sphere2500 dataset, an exception to this trend is observed when the range measurement probability is 1.0. Specifically, the eight-agent system demonstrates improved convergence and solution precision compared to the four-agent system. We conjecture that the increase in range measurement density compensates for the loss in connectivity in each agent’s underlying pose graph, as intra-agent relative pose loop closures are redistributed as inter-agent loop closures. Observations on SDP tightness further support this conjecture, as increasing range measurement density improves tightness as the number of agents increases.
V Conclusion
In this work, we presented DCORA, the first distributed certifiably correct algorithm for range-aided simultaneous localization and mapping (RA-SLAM). DCORA generalizes the Riemannian Staircase method established for distributed certifiably correct pose graph optimization to the RA-SLAM problem and demonstrates comparable solution quality to its centralized counterpart, CORA, on real-world datasets. The structure of the RA-SLAM problem in the distributed setting was studied using parametrically varied synthetic datasets, revealing the number of agents to be the predominant factor affecting DCORA’s convergence rate and solution precision. Results also showed that graph connectivity influences SDP tightness, where additional range measurements can aid solution quality for systems distributing relative pose measurements across a large number of agents. In future work, we aim to expand the real-world evaluation of DCORA across larger multi-agent systems under various noise regimes. We also plan on studying DCORA’s runtime performance with a fully distributed implementation. Of broader importance to DCORA’s method class, efforts to significantly enhance the speed, scalability, and communication efficiency of Riemannian block coordinate descent by incorporating (at least partial) second-order information into its design would be a substantial contribution towards enabling efficient estimation and control methods for multi-agent systems in general. This, in addition to identifying sufficient conditions to ensure that DCORA’s (and CORA’s) SDP relaxation remains exact under bounded noise for the general RA-SLAM problem, are directions for future work.